OpenStudy (anonymous):

simplify cos^2 Ө / 1 - sin Ө

OpenStudy (anonymous):

cos^2 = 1 - sin^2

OpenStudy (anonymous):

none of my answer choices have cos in them

OpenStudy (anonymous):

no, replace cos^2 with 1 - sin^2. Then simplify

OpenStudy (anonymous):

it's like @10115658 explained use Cos^2θ =1-Sin^2θ

OpenStudy (anonymous):

there isn't an " ^2 " option either

OpenStudy (anonymous):

these are the answer choices: sin Ө 1 + sin Ө 1 - sin Ө 1 - sin Ө / sin Ө

OpenStudy (anonymous):

\[\cos(\Theta)^2 = 1 - \sin(\Theta)^2 = (1-\sin(\Theta))(1+\sin(\Theta))\] then : \[\frac{ \cos(\Theta )^2}{ 1-\sin(\Theta) } = \frac{ (1-\sin(\Theta)(1+\sin(\Theta)) }{ 1-\sin(\Theta) } = ?\] Let me know if you got it :)

OpenStudy (anonymous):

What should cancel?

OpenStudy (anonymous):

the 1 - sin

OpenStudy (anonymous):

leaving what

OpenStudy (anonymous):

1 + sin Ө

OpenStudy (anonymous):

So what's the answer

OpenStudy (anonymous):

1 + sin Ө

OpenStudy (anonymous):

correct A

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

great explanation @Omar_Elboredy

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

Wlc @TheForbiddenFollower @10115658 :)